# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ E.2.2 Additional results
Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) Top-$k$ sampling 0.635 0.814 0.113 0.585 0.073 0.214 0.400 0.590 0.486 0.176 0.315
LoRA-MLE ($r=8$) Typical $p$ sampling 0.645 0.817 0.108 0.589 0.077 0.215 0.402 0.587 0.492 0.177 0.319
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.646 0.818 0.106 0.592 0.074 0.214 0.401 0.589 0.489 0.178 0.318
LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.642 0.816 0.110 0.600 0.079 0.214 0.406 0.588 0.500 0.175 0.322
LoRA-MCL ($\varepsilon=0.0005$) Typical $p$ sampling 0.648 0.815 0.111 0.595 0.082 0.215 0.409 0.584 0.506 0.175 0.327
LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.649 0.817 0.107 0.595 0.079 0.214 0.408 0.586 0.504 0.171 0.325
Table 8: Results for AudioCaps with 5 candidates and Sampling-based decoding.
Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) Top-$k$ sampling 0.562 0.810 0.134 0.560 0.115 0.299 0.457 0.699 0.442 0.241 0.318
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.576 0.810 0.141 0.560 0.121 0.299 0.460 0.699 0.450 0.241 0.325
LoRA-MLE ($r=8$) Typical sampling 0.567 0.809 0.143 0.550 0.115 0.300 0.452 0.697 0.390 0.241 0.294
LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.583 0.830 0.098 0.582 0.115 0.298 0.470 0.700 0.498 0.239 0.344
LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
LoRA-MCL ($\varepsilon=0.0005$) Typical sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
E.3 Image Captioning Experiments
E.3.1 Experimental setup
Table 9: Quality and Diversity Evaluation on TextCaps with $3$ candidates. For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ). LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=96$. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
@@ E.3.1 Experimental setup
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE Beam Search 3 0.433 0.509 0.688 0.802 0.318 0.315 0.580 0.670 1.517 0.244 0.873
LoRA-MLE Beam Search 3 0.383 0.421 0.833 0.788 0.315 0.317 0.573 0.670 1.517 0.241 0.874
LoRA-MLE Beam Search 6 0.402 0.457 0.786 0.795 0.338 0.326 0.583 0.671 1.557 0.246 0.895
LoRA-MLE Beam Search 6 0.401 0.456 0.784 0.796 0.339 0.327 0.585 0.672 1.572 0.248 0.903
LoRA-MLE DBS ($\lambda=0.5$) 3 0.489 0.600 0.529 0.818 0.345 0.325 0.596 0.684 1.571 0.253 0.902
LoRA-MLE DBS ($\lambda=0.8$) 3 0.530 0.655 0.437 0.824 0.349 0.327 0.601 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.540 0.669 0.416 0.822 0.348 0.326 0.599 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=0.5$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=0.8$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=1.0$) 3 0.535 0.665 0.425 0.827 0.357 0.327 0.601 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.5$) 6 0.455 0.532 0.694 0.813 0.344 0.328 0.595 0.681 1.580 0.253 0.908
LoRA-MLE DBS ($\lambda=0.8$) 6 0.470 0.549 0.671 0.812 0.341 0.328 0.593 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.474 0.553 0.666 0.812 0.340 0.328 0.592 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=0.5$) 6 0.451 0.528 0.697 0.808 0.343 0.329 0.594 0.681 1.586 0.253 0.911
LoRA-MLE DBS ($\lambda=0.8$) 6 0.464 0.542 0.681 0.810 0.340 0.329 0.593 0.680 1.583 0.253 0.909
LoRA-MLE DBS ($\lambda=1.0$) 6 0.473 0.551 0.671 0.810 0.341 0.328 0.592 0.680 1.584 0.251 0.908
LoRA-MCL Beam Search 1 0.499 0.599 0.520 0.828 0.344 0.330 0.597 0.690 1.674 0.255 0.955
LoRA-MCL Beam Search 2 0.507 0.618 0.490 0.824 0.360 0.333 0.604 0.687 1.627 0.258 0.932
@@
We used LLaVA 1.6 as the base model, which features 7.1 billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$, up and down linear projections of each block of the language model. We used a rank $r$, with $r=8$ and $\alpha=32$ unless otherwise stated, with dropout equal to $0.1$. We trained each model with a batch size of $8$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with maximum learning rate of $10^{-4}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. We used 1 epoch for training, where we duplicated the image as many times as the number of its captions, such that the model sees exactly one time each caption.
E.4 Additional results

### v2

@@ E.6.4 Additional results
Training Decoding $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
LoRA-MLE ($r=8$) Top-$k$ sampling 0.813 0.109 0.066 0.215 0.402 0.593 0.497 0.176 0.323
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.812 0.112 0.073 0.212 0.403 0.586 0.516 0.173 0.330
LoRA-MLE ($r=8$) Typical $p$ sampling 0.812 0.111 0.073 0.212 0.403 0.586 0.516 0.173 0.331
LoRA-MCL (annealed) Top-$k$ sampling 0.819 0.097 0.074 0.212 0.403 0.597 0.507 0.175 0.327
LoRA-MCL (annealed) Nucleus (Top-$p$) sampling 0.820 0.098 0.074 0.212 0.403 0.589 0.507 0.171 0.325
LoRA-MCL (annealed) Typical $p$ sampling 0.819 0.103 0.072 0.211 0.402 0.590 0.509 0.172 0.327
Table 10: Results for AudioCaps with 5 hypotheses and MAP Decoding.
@@
Training Decoding Beam $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
LoRA-MLE ($r=8$) BS 5 0.395 0.764 0.267 0.377 0.606 0.700 1.121 0.250 0.668
LoRA-MLE ($r=40$) BS 5 0.392 0.773 0.280 0.382 0.606 0.701 1.144 0.251 0.681
LoRA-MLE ($r=8$) BS 10 0.410 0.746 0.286 0.385 0.610 0.704 1.157 0.256 0.689
LoRA-MLE ($r=40$) BS 10 0.407 0.747 0.298 0.382 0.610 0.704 1.151 0.255 0.686
LoRA-MLE ($r=8$) BS 25 0.417 0.732 0.281 0.383 0.609 0.706 1.144 0.258 0.683
LoRA-MLE ($r=40$) BS 25 0.415 0.735 0.289 0.384 0.611 0.706 1.152 0.260 0.688
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.553 0.448 0.268 0.378 0.610 0.708 1.117 0.248 0.662
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.557 0.444 0.263 0.380 0.612 0.707 1.130 0.248 0.669
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.580 0.403 0.275 0.376 0.612 0.708 1.128 0.249 0.667
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.580 0.403 0.268 0.378 0.614 0.709 1.138 0.250 0.672
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.481 0.684 0.274 0.373 0.606 0.709 1.114 0.259 0.665
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.481 0.683 0.278 0.380 0.610 0.710 1.124 0.257 0.670
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.486 0.678 0.276 0.372 0.607 0.710 1.115 0.259 0.665
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.492 0.675 0.276 0.376 0.609 0.709 1.118 0.256 0.666
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.402 0.756 0.273 0.376 0.607 0.703 1.129 0.251 0.673
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.399 0.763 0.282 0.381 0.605 0.702 1.139 0.253 0.680
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.402 0.757 0.271 0.377 0.608 0.703 1.128 0.251 0.672
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.398 0.765 0.282 0.381 0.606 0.701 1.136 0.252 0.678
LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 1 0.385 0.731 0.198 0.343 0.574 0.693 0.969 0.226 0.584
LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 1 0.479 0.588 0.207 0.352 0.586 0.695 0.992 0.231 0.597
LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 1 0.576 0.438 0.185 0.332 0.567 0.683 0.916 0.218 0.550
LoRA-MLE ($r=8$) TTA BS ($\phi_{1}$) 5 0.395 0.733 0.241 0.358 0.592 0.700 1.057 0.242 0.636
LoRA-MLE ($r=8$) TTA BS ($\phi_{2}$) 5 0.491 0.590 0.260 0.360 0.597 0.703 1.047 0.246 0.630
LoRA-MLE ($r=8$) TTA BS ($\phi_{3}$) 5 0.596 0.435 0.251 0.347 0.586 0.693 1.005 0.235 0.605
LoRA-MoE BS 5 0.396 0.766 0.274 0.381 0.608 0.702 1.129 0.252 0.674
LoRA-MoE BS 10 0.411 0.746 0.288 0.385 0.613 0.703 1.161 0.256 0.692
LoRA-MoE BS 25 0.415 0.736 0.287 0.385 0.612 0.705 1.154 0.258 0.689
LoRA-MoE DBS ($\lambda=0.8$) 5 0.555 0.443 0.275 0.379 0.611 0.707 1.136 0.248 0.670
LoRA-MoE DBS ($\lambda=1.0$) 5 0.578 0.409 0.268 0.377 0.609 0.708 1.130 0.247 0.667
LoRA-MoE DBS ($\lambda=0.8$) 10 0.484 0.677 0.279 0.374 0.608 0.709 1.122 0.256 0.667
LoRA-MoE DBS ($\lambda=1.0$) 10 0.488 0.675 0.283 0.374 0.608 0.710 1.119 0.257 0.666
LoRA-MoE DBS ($\lambda=0.8$) 25 0.402 0.757 0.275 0.382 0.609 0.701 1.135 0.253 0.676
LoRA-MoE DBS ($\lambda=1.0$) 25 0.401 0.759 0.275 0.382 0.609 0.701 1.133 0.253 0.675
LoRA-MoE SR BS 1 0.256 0.909 0.165 0.318 0.538 0.670 0.865 0.200 0.526
LoRA-MoE SR BS 2 0.269 0.895 0.195 0.326 0.549 0.675 0.905 0.210 0.551
LoRA-MoE SR BS 5 0.261 0.907 0.196 0.331 0.550 0.678 0.925 0.216 0.563
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.580 0.374 0.259 0.377 0.613 0.711 1.119 0.248 0.662
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.551 0.427 0.273 0.382 0.622 0.712 1.181 0.253 0.695
LoRA-MCL (annealed) BS 1 0.570 0.397 0.275 0.379 0.615 0.713 1.149 0.255 0.680
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.575 0.401 0.277 0.385 0.626 0.712 1.152 0.258 0.684
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.544 0.464 0.298 0.394 0.631 0.712 1.209 0.260 0.714
LoRA-MCL (annealed) BS 2 0.574 0.411 0.309 0.392 0.632 0.714 1.205 0.264 0.712
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.579 0.410 0.306 0.392 0.631 0.712 1.190 0.263 0.706
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.542 0.491 0.309 0.399 0.636 0.715 1.237 0.265 0.728
LoRA-MCL (annealed) BS 5 0.575 0.423 0.315 0.398 0.636 0.714 1.211 0.268 0.716
Table 11: Results for AudioCaps with 5 hypotheses and Sampling-based Decoding.
Training Decoding $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
LoRA-MLE ($r=8$) Top-$k$ sampling 0.711 0.173 0.169 0.335 0.549 0.701 0.906 0.229 0.543
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.697 0.193 0.180 0.344 0.564 0.704 0.940 0.234 0.563
LoRA-MLE ($r=8$) Typical $p$ sampling 0.699 0.189 0.182 0.343 0.564 0.704 0.945 0.233 0.566
LoRA-MCL (annealed) Top-$k$ sampling 0.735 0.149 0.175 0.342 0.568 0.705 0.938 0.233 0.563
LoRA-MCL (annealed) Nucleus (Top-$p$) sampling 0.724 0.163 0.182 0.340 0.566 0.702 0.950 0.235 0.569
LoRA-MCL (annealed) Typical $p$ sampling 0.724 0.160 0.186 0.341 0.568 0.703 0.953 0.235 0.570
E.6.5 Qualitative Examples
We provide some qualitative examples of the predictions on AudioCaps. Here, LoRA-MCL uses $\varepsilon=0.05$, Beam Search with $B=5$ and $K=5$, LoRA-MoE uses Beam search with $B=25$.
Example 1. References:
  • A large truck driving by as an emergency siren wails and truck horn honks
  • A wailing siren fades, a motor sputters, then the siren resumes accompanied by blaring horns
  • An emergency siren ringing with car horn honking
  • A fire truck engine runs and the siren is blowing but stops, traffic is present, the fire truck horn honks twice, and the siren begins again
  • A fire engine with a siren fading then another loud siren
LoRA-MoE.
An emergency vehicle passes by and blows its horn followed by the siren of an emergency vehicle}\}
[…]
LoRA-MCL.
An emergency vehicle siren is going off}\}
[…]
Example 2. References:
  • A man speaks as birds chirp and dogs bark
  • A man is speaking as birds are squawking, and a dog barks
  • A man talks while several animals make noises in the background
  • A man speaking followed by dogs barking alongside chimps screaming and birds chirping
  • A man speaking as monkeys scream and dogs bark followed by birds cawing in the distance
LoRA-MoE.
A man is speaking and dogs are barking}\}
[…]
LoRA-MCL.
A man speaking and dogs barking}\}
[…]
Example 3. References:
  • An engine running and wind with various speech in the background
  • A motorboat engine operating as a crowd of people talk followed by metal creaking and a man speaking
  • A large motor is running smoothly, water is splashing, people are talking in the background, and an adult male speaks in the distance
  • A ship engine running as a crowd of people talk followed by a ship hull creaking as wind blows heavily into a microphone
  • Outdoor noise from a water vehicle as people are talking
LoRA-MoE.
Humming of an idling engine followed by a horn honking}\}
[…]
LoRA-MCL.
A boat motor is running and people are talking in the background}\}
[…]
E.7 Image Captioning Experiments
E.7.1 Experimental setup
Table 12: Quality and Diversity Evaluation on TextCaps with 33 candidates. For each of the presented metrics, higher is better $(\uparrow)$ except for mBLEU-4 $(\downarrow)$. LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=96$. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with . PaliGemma-3B rows are shown in gray, as this model was fully fine-tuned and is therefore not directly comparable to LoRA-based methods.
Training Decoding Beam $\mathrm{Div}_{2}$ $\mathrm{mBLEU}_{4}$ $\mathrm{BLEU}_{1}$ $\mathrm{BLEU}_{4}$ $\mathrm{METEOR}$ $\mathrm{ROUGE}_{L}$ $\mathrm{sBERT}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
LoRA-MLE BS 3 0.509 0.688 0.802 0.318 0.315 0.580 0.670 1.517 0.244 0.873
LoRA-MLE BS 6 0.457 0.786 0.795 0.338 0.326 0.583 0.671 1.557 0.246 0.895
LoRA-MLE BS 3 0.421 0.833 0.788 0.315 0.317 0.573 0.670 1.517 0.241 0.874
LoRA-MLE BS 6 0.456 0.784 0.796 0.339 0.327 0.585 0.672 1.572 0.248 0.903
LoRA-MLE DBS ($\lambda=0.5$) 3 0.600 0.529 0.818 0.345 0.325 0.596 0.684 1.571 0.253 0.902
LoRA-MLE DBS ($\lambda=0.8$) 3 0.655 0.437 0.824 0.349 0.327 0.601 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.669 0.416 0.822 0.348 0.326 0.599 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=0.5$) 6 0.532 0.694 0.813 0.344 0.328 0.595 0.681 1.580 0.253 0.908
LoRA-MLE DBS ($\lambda=0.8$) 6 0.549 0.671 0.812 0.341 0.328 0.593 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.553 0.666 0.812 0.340 0.328 0.592 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=0.5$) 3 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=0.8$) 3 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=1.0$) 3 0.665 0.425 0.827 0.357 0.327 0.601 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.5$) 6 0.528 0.697 0.808 0.343 0.329 0.594 0.681 1.586 0.253 0.911
LoRA-MLE DBS ($\lambda=0.8$) 6 0.542 0.681 0.810 0.340 0.329 0.593 0.680 1.583 0.253 0.909
LoRA-MLE DBS ($\lambda=1.0$) 6 0.551 0.671 0.810 0.341 0.328 0.592 0.680 1.584 0.251 0.908
LoRA-MoE DBS ($\lambda=0.5$) 3 0.600 0.529 0.822 0.347 0.327 0.598 0.684 1.610 0.258 0.923
LoRA-MoE DBS ($\lambda=0.8$) 3 0.654 0.441 0.828 0.353 0.327 0.603 0.685 1.616 0.254 0.924
LoRA-MoE DBS ($\lambda=1.0$) 3 0.666 0.421 0.828 0.354 0.328 0.602 0.685 1.622 0.253 0.926
LoRA-MoE DBS ($\lambda=0.5$) 6 0.530 0.698 0.814 0.348 0.331 0.595 0.681 1.607 0.257 0.923
LoRA-MoE DBS ($\lambda=0.8$) 6 0.545 0.678 0.813 0.349 0.330 0.596 0.680 1.608 0.255 0.922
LoRA-MoE DBS ($\lambda=1.0$) 6 0.551 0.670 0.813 0.346 0.330 0.596 0.679 1.607 0.254 0.921
LoRA-MoE SR BS 1 0.556 0.597 0.809 0.315 0.311 0.576 0.678 1.541 0.243 0.883
PaliGemma-3B (ft) DBS ($\lambda=0.5$) 3 0.595 0.532 0.832 0.369 0.330 0.608 0.688 1.658 0.255 0.947
PaliGemma-3B (ft) DBS ($\lambda=0.8$) 3 0.635 0.467 0.835 0.374 0.332 0.611 0.689 1.661 0.257 0.949
PaliGemma-3B (ft) DBS ($\lambda=1.0$) 3 0.655 0.439 0.833 0.368 0.332 0.609 0.687 1.654 0.256 0.944
LoRA-MCL BS 1 0.599 0.520 0.828 0.344 0.330 0.597 0.690 1.674 0.255 0.955
LoRA-MCL BS 2 0.618 0.490 0.824 0.360 0.333 0.604 0.687 1.627 0.258 0.932
@@ E.7.1 Experimental setup
We used LLaVA 1.6 with Vicuna-7B (Zheng et al., 2023) for the LLM, as the base model, which features 7.1 billion parameters. We used the official codebase for the implementation. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$, up and down linear projections of each block of the language model. Fine-tuning hyperparameters follow those from the recipe provided by the authors, except that we used smaller batch sizes and a lower LoRA rank due to compute constraints. We used a rank $r$, with $r=8$ and $\alpha=32$ unless otherwise stated, with dropout equal to $0.1$ (enabled during training, and disabled during inference). We trained each model with a batch size of $8$, with Adam optimizer (Diederik and Jimmy, 2014) (with $\beta_{1}=0.9$, and $\beta_{2}=0.999$), using a cosine scheduler with maximum learning rate of $2\times 10^{-4}$, with a warmup ratio of $0.03$. Maximum sequence length is set to $2048$. Gradient clipping is used with a maximum gradient norm of $1.0$. We used 1 epoch for training, where we duplicated the image as many times as the number of its captions, such that the model sees exactly one time each caption.
E.7.2 Additional results
Comparing with PaliGemma-3B. Several works have explored improving caption diversity in image captioning (Wang and Chan, 2019; Wang et al., 2020; Mahajan and Roth, 2020). Among those evaluating on TextCaps, Zhang et al. (2022) and Xu et al. (2021) report results, but their performance is far below that of recent VLMs, such as PaliGemma-3B (Beyer et al., 2024). For example, the reported best corpus-level CIDEr on the TextCaps validation set are: 76.6 for Zhang et al. (2022), 95.5 for Xu et al. (2021), 127.48 for PaliGemma-3B (224x224). We therefore compare with PaliGemma-3B, which provides a publicly available fine-tuned version on Hugging Face. We evaluate it using DBS returning $K=3$ hypotheses (Table 12). PaliGemma is a strong baseline, arguably the state-of-the-art open-weight model on TextCaps. Note, however, that the comparison is not entirely fair, as PaliGemma undergoes full-weight fine-tuning (instead of LoRA). Despite this, our method matches and even slightly improves its performance (Oracle SPIDEr of $0.955$ vs. $0.949$). We also expect that applying LoRA-MCL to PaliGemma could further improve the results.